OpenStudy (anonymous):

Laurie inherited 40% of her father's estate. After paying a tax equal to 30% of her inheritance, what percent of her father's estate did she own?

OpenStudy (ganpat):

lets the fathers state be x ... and she inherited 40 % i.e. = 40% 0f x Now she pays an tax of 30% on that.. So what left with her is 100- 30 = 70% of what she got, that is equal to... 70% of 40% of x = (70/ 100) * (40/100) * x (28/ 100 ) * x = 28% 0f x...